OpenStudy (anonymous):

how do i simplify (1-(x/(1+x)))/(1+(x/(1+x)))

OpenStudy (empty):

It's going to be sorta difficult, but the trick is to multiply by a clever form of the number 1, since multiplying by one doesn't change anything. Here's how it works, and you can use it anywhere: This is my own made up example, so I used n instead of x so I don't confuse you. \[\frac{1}{1+\frac{1}{n}} =\frac{1}{(1+\frac{1}{n})} \frac{n}{n} = \frac{n}{n+1} \] It's sorta tricky so you might have to think about it for a bit and if you're stuck I can help you out.

OpenStudy (anonymous):

yes i understand the concept:D i'm just having trouble with this specific example the double division is throwing me off D:

OpenStudy (empty):

try something out, the worst that could happen is that you accidentally get it right. Then I can see how to help you out, like what part's confusing.

OpenStudy (anonymous):

okay one second ill try and type it all out

OpenStudy (empty):

if you want it might be easier to draw it out

OpenStudy (anonymous):

okay well i have like a page of trial and error ill just tell you what im doing wrong other wise itll take me forever D:

OpenStudy (anonymous):

okay so i isolated the numerator and the denominator and gave the numerator a common denominator of (1+(x/1+x))

OpenStudy (anonymous):

wait no thats stupid im multiplying i dont need a common denominator .... ugh half a bloody page of useless work

OpenStudy (empty):

Hey well that's a good catch though! You probably won't make that mistake again on a test where it actually counts at least.

OpenStudy (empty):

Just to be sure before I give you a hint, is this the question: \[\Large \frac{1-\frac{x}{1+x}}{1+\frac{x}{1+x}}\] I just want to make sure that's what you're looking at cause there's a lot of stuff in there that can mess up haha.

OpenStudy (usukidoll):

complex fractions.

OpenStudy (anonymous):

you know what it'd be rather hectic explaining this through text or using that retched drawing tool, i think iv got it from here. Go help Some one else ill just use a calculator or something:)

OpenStudy (empty):

Well alright haha I'll give you a hint anyways. Multiply by this form of 1 and distribute it out: \[\frac{1+x}{1+x}\] good luck

OpenStudy (anonymous):

thanks !